# On-Chip Frequency Noise Cancellation in Nanomechanical Resonators using Cavity Optomechanics

### v1

@@ IV Mass sensing simulated experiment
We assign the improved SNR of the signal at $\omega_{\mathrm{d}}$ relative to that measured at $\omega_{1}$ to the cancellation of intrinsic frequency fluctuations at low $\Omega$. As we see in Fig. 2(h), this effect is significant up to roughly $\Omega=$1\text{\,}\mathrm{Hz}$\$, explaining the reduced frequency fluctuations and the improved SNR in Fig. 3. However, for $\Omega>$1\text{\,}\mathrm{Hz}$\$, we can see in Fig. 2(h) that the signal at $\omega_{\mathrm{d}}$ has a higher $S_{\omega}$ than $\omega_{1}$, due to the increased thermomechanical force noise that it inherits from $\omega_{2}$. This noise contribution can be mitigated by averaging the data over time. This is why the SNR at $\omega_{\mathrm{d}}$ increases from 4.1 to 6.2 when applying the moving-average filter, while the one at $\omega_{1}$ remains the same within the measurement uncertainty. In principle, the low-$\Omega$ fluctuations can also be reduced in post-processing by calculating the signal $\omega^{\mathrm{cal}}_{\mathrm{d}}=\omega_{2}-\omega_{1}$ [30]. However, the SNR of this signal always remains below that of $\omega_{\mathrm{d}}$ because it samples the detection noise twice, once when measuring $\omega_{1}$ and once for $\omega_{2}$. Assuming that the detection noise imprinted on the two PLL measurements is uncorrelated, the resulting high-$\Omega$ noise present for $\omega^{\mathrm{cal}}_{\mathrm{d}}$ should be roughly $\sqrt{2}$ times larger than that for $\omega_{\mathrm{d}}$. This is in agreement with the observation that averaging improves the SNR at $\omega^{\mathrm{cal}}_{\mathrm{d}}$ significantly. Nevertheless, it never quite catches up with the on-chip noise cancellation we demonstrate with $\omega_{\mathrm{d}}$.
V Conclusions and outlook
@@
In this work, we perform frequency noise characterization of high-$Q$ devices and observe correlated intrinsic frequency fluctuations of different mechanical modes. This correlation for low mass [17], high-$Q$ resonators is expected to predominantly be correlated flicker [15], resulting from the shared defects interacting with the various mechanical modes. Using nonlinear optomechanical transduction, we show how the correlation can be used for on-chip frequency noise reduction. With this method, we achieve a 20 fold decrease in absolute Allan Deviation at timescales larger than 1 second. Furthermore, the method performs a factor $3.2$ better than off-chip frequency noise reduction by post-processing, attributed to the fact that the method is not limited by PLL errors in the frequency measurements of individual modes, as the detection noise is only sampled once, not independently for both modes. Finally, we verify the usability of this noise suppression method in a proof-of-principle experiment that mimics the conditions for mass sensing. We find similar behaviour across multiple devices and by engaging different modes. Therefore, we conclude that our method can be a useful resource for precise frequency sensing experiments, such as mass spectroscopy [7, 8] or photothermal particle detection [10]. High-precision frequency tracking, combined with the excellent force sensitivity and low detection noise of these perimeter-mode resonators, is also crucial for nuclear spin detection [44, 45], and ultimately for nanoscale magnetic resonance imaging [46].
@@
Acknowledgements.
This research has been possible thanks to the Swiss National Science Foundation (SNSF) through grant no. 200021_200412 and the Novo Nordisk Foundation through grant no. NNF22OC0077964. N.J.E. acknowledges funding from the European Research Council (ERC) under grant no. 101117144. We would like to thank Prof Albert Schliesser (NBI Copenhagen), Prof Silvan Schmid (TU Vienna), Hajrudin Bešić (TU Vienna) and Olivier Faist (Zurich Instruments) for insightful discussions and nudges; Diego Visani, Thomas Gisler, Nils Prumbaum, Urs Grob and Luis Mestre for in-lab assistance. The technical support for the vacuum-compatible fibre alignment stage from Walter Bachmann of the engineering office at Dept. of Physics, ETH Zürich is highly appreciated.
@@ Appendix B Optomechanical Characterisation
Figure 6: Ringdown measurements of the perimeter resonator’s modes $\omega_{1}$ in red and $\omega_{2}$ in blue with their respective fits (dashed).
@@ Appendix C Optical frequency noise calibration
where
$D^{(0)}$ $=1-\frac{4\eta(1-\eta)}{1+\bar{\delta}^{2}},$ (1)
$D^{(1)}$ $=4\eta(1-\eta)\frac{2\bar{\delta}}{(1+\bar{\delta}^{2})^{2}},$ (2)
and $\xi(t)$ is the cavity frequency noise normalized to $\kappa/2$, containing the mechanical displacement. See later in SI for the derivations. The technique used here is based on using the to use the DC component of the photocurrent to calibrate the AC component. The DC and AC components of the detected voltage is given by
$V_{\mathrm{DC}}$ $=Z_{\mathrm{DC}}D^{(0)},$ (3)
$V_{\mathrm{AC}}(t)$ $=Z_{\mathrm{AC}}D^{(1)}\xi(t),$ (4)
where $Z_{\mathrm{DC}}$ and $Z_{\mathrm{AC}}$ are values of the trans-impedance gain at DC and AC frequencies. For the power spectral density (PSD) of the AC signal, defined as $S_{VV}(\omega)=\int e^{i\omega\tau}\langle V_{\mathrm{AC}}(t)V_{\mathrm{AC}}(t+\tau)\rangle d\tau$, it can be written as
@@
This conversion is used to calibrate the data presented in Fig. 1(e). To calibrate the vacuum optomechanical coupling rate $g_{0}$ one has to integrate both side of Eq. 5 around a mechanical sideband. Integration of $S_{\xi\xi}$ around a mechanical mode with frequency $\omega_{m}$ and OM coupling $g_{0}$ obtains $8g_{0}^{2}n_{\mathrm{th}}/k^{2}$ where $n_{\mathrm{th}}=k_{B}T/\hbar\omega$ is the thermal occupation of the mode. Using this, $g_{0}$ is given by
@@ Appendix D Frequency noise characterization
When measuring the mechanical frequency fluctuations , it is known that the measurement is contaminated by three main sources of noise: thermomechanical, detection, and intrinsic frequency noise. Together, they constitute the total added frequency noise, whose power spectral density is given by:
$S^{\text{tot}}_{\omega}(\Omega)=S_{\omega}^{\mathrm{th}}(\Omega)+S_{\omega}^{\mathrm{det}}(\Omega)+S_{\omega}^{\mathrm{int}}(\Omega).$ (7)
It is known that the thermomechanical term has no frequency dependence and that the detection term has a quadratic frequency scaling. For the data shown in the main-text Fig. 1(f), we use the model $S_{\omega}(\Omega)=B\Omega^{2}+A/\Omega$ for fitting and extracting the $1/\Omega$ noise component. We have not included a constant term since as mentioned in the main text, the thermomechanical component for the high-$Q$ OOP mode is too low to beresolved. In Fig. 7 we show an example of the full f-PSD, the segment used for fitting and the result of the fit for the lowest input optical power. The roll-off at frequencies above $1\text{\,}\mathrm{kHz}$ is a result of the finite bandwidth of the lock-in amplifier’s demodulator. We do not include this segment in the model. The initial points of the data are not included either due to the lack of a sufficient number of points and large variations. $1/\Omega$-type frequency noise with a frequency noise PSD $S_{\omega}(\Omega)=A/\Omega$ corresponds to the constant Allan variance with standard deviation $\sigma^{2}_{\omega}(\tau)=2A\ln{(2)}$. This relation is used to produce the inset of the main-text Fig. 1(f).
@@
We also observe the influence of driving power and optical power on the overall frequency noise. We observe that while both parameters influence the detector branch, the effect is differently observed. Tuning the drive power affects the overall signal strength while tuning the optical power reduces the overall noise floor, as we can correlate from the side peak around $1.2\text{\,}\mathrm{kHz}$ for the $\omega_{1}$ mode, which remains the same height above noise for changing drive power but becomes more pronounced with increasing optical power. On the second hand, decreasing optical power does not influence the thermomechanical noise floor but a decreasing drive power causes an increase in thermomechanical noise for the low-$Q$ mode at $\omega_{2}$.
@@
Appendix E Optical Non-linear Transduction for difference Frequency generation

### v2

@@ Abstract
We assign the increased SNR of the signal at $\omega_{\mathrm{d}}$ relative to that measured at $\omega_{1}$ to the cancellation of intrinsic frequency fluctuations at low $\Omega$. As we see in Fig. 2(h), this effect is significant up to roughly $\Omega=$1\text{\,}\mathrm{Hz}$\$, explaining the reduced frequency fluctuations and the improved SNR in Fig. On-Chip Frequency Noise Cancellation in Nanomechanical Resonators using Cavity Optomechanics. However, for $\Omega>$1\text{\,}\mathrm{Hz}$\$, we can see in Fig. 2(h) that the signal at $\omega_{\mathrm{d}}$ has a greater $S_{\omega}$ than $\omega_{1}$ due to the increased thermomechanical force noise that it inherits from $\omega_{2}$. This noise contribution can be mitigated by averaging the data over time. This is why the SNR at $\omega_{\mathrm{d}}$ increases from 4.1 to 6.2 when the moving-average filter is applied, while the one at $\omega_{1}$ remains constant within the measurement uncertainty. In principle, the low-$\Omega$ fluctuations can also be reduced in postprocessing by calculating the signal $\omega^{\mathrm{cal}}_{\mathrm{d}}=\omega_{2}-\omega_{1}$ [32]. However, the SNR of this signal always remains below that of $\omega_{\mathrm{d}}$ because it samples the detection noise twice, once when $\omega_{1}$ is measured and once for $\omega_{2}$. Assuming that the detection noise imprinted on the two PLL measurements is uncorrelated, the resulting high-$\Omega$ noise present for $\omega^{\mathrm{cal}}_{\mathrm{d}}$ should be roughly $\sqrt{2}$ times larger than that for $\omega_{\mathrm{d}}$. This is in agreement with the observation that averaging increases the SNR at $\omega^{\mathrm{cal}}_{\mathrm{d}}$ significantly. Nevertheless, it never quite catches up with the on-chip noise cancellation we demonstrate with $\omega_{\mathrm{d}}$.
@@
Conclusions and outlook.— In this work, we perform frequency noise characterization of high-$Q$ devices and observe correlated intrinsic frequency fluctuations of different mechanical modes. This correlation for low-mass [19], high-$Q$ resonators is expected to predominantly be correlated $1/\Omega$-noise [15], resulting from the shared defects interacting with the various mechanical modes. Using nonlinear optomechanical transduction, we show how the correlation can be used for on-chip frequency noise reduction. With this method, we achieve a 20-fold decrease in absolute Allan deviation at times larger than $1\text{\,}\mathrm{s}$. Furthermore, from comparison of the SNRs, the method performs up to a factor $3.2$ better than off-chip frequency noise reduction by postprocessing, which is attributed to the method not being limited by PLL errors in the frequency measurements of individual modes, as the detection noise is sampled only once, not independently for both modes. Finally, we verify the usability of this noise suppression method in a proof-of-principle experiment that mimics the conditions for mass sensing. We find similar behavior across multiple devices and by engaging different modes. Therefore, we conclude that our method can be a useful resource for precise frequency sensing experiments, such as mass spectroscopy [7, 8] or photothermal particle detection [10]. High-precision frequency tracking, combined with the excellent force sensitivity and low detection noise of these perimeter-mode resonators, is also crucial for nuclear spin detection [50, 51], and ultimately for nanoscale magnetic resonance imaging [52].
@@
Acknowledgments. This research was possible thanks to the Swiss National Science Foundation (SNSF) through Grants No. 200021_200412 and No. 200021E_209235, and the Novo Nordisk Foundation through Grant No. NNF22OC0077964. N.J.E. acknowledges funding from the European Research Council (ERC) under Grant No. 101117144. We would like to thank Albert Schliesser (Niels Bohr Institute, University of Copenhagen), Silvan Schmid (TU Vienna), Hajrudin Bešić (TU Vienna), and Olivier Faist (Zurich Instruments) for insightful discussions and nudges; Diego Visani, Thomas Gisler, Nils Prumbaum, Urs Grob and Luis Mestre are acknowledged for in-laboratory assistance. The technical support from Walter Bachmann of the engineering office and from the mechanics workshop of the Department of Physics at ETH Zürich, is highly appreciated.
Data Availability.— The data that support the findings of this article are openly available. [54].
Supplemental Material for:
On-Chip Frequency Noise Cancellation in Nanomechanical Resonators
using Cavity Optomechanics
@@ Appendix B Optomechanical Characterisation
Figure 4: Optomechanical coupling. (a) A broadband spectrum obtained in direct detection from the device. The red line corresponds to the thermomechanical spectrum acquired when the laser is red detuned from the cavity. The peaks correspond to the thermal motion of different modes of the polygon resonator, a few of which are marked alongside their mode shapes. (b) FEM simulation of the electric field intensity
@@
Figure 5: Ringdown measurements of the device’s perimeter modes $\omega_{1}$ in red and $\omega_{2}$ in blue with their respective fits (dashed).
@@ Appendix C Optical frequency noise calibration
where {align} D^(0) = 1 - 4η(1-η)1 + ¯δ2,
D^(1) = 4η(1-η)2¯δ(1 + ¯δ2)2, and $\xi(t)$ is the cavity frequency noise normalized to $\kappa/2$, containing the mechanical displacement. See later in SI for the derivations. The technique used here is based on using the to use the DC component of the photocurrent to calibrate the AC component. The DC and AC components of the detected voltage is given by {align} V_DC = Z_DCD^(0),
V_AC(t) = Z_ACD^(1)ξ(t), where $Z_{\mathrm{DC}}$ and $Z_{\mathrm{AC}}$ are values of the trans-impedance gain at DC and AC frequencies. For the power spectral density (PSD) of the AC signal, defined as $S_{VV}(\omega)=\int e^{i\omega\tau}\langle V_{\mathrm{AC}}(t)V_{\mathrm{AC}}(t+\tau)\rangle d\tau$, it can be written as
@@
This conversion is used to calibrate the data presented in the main-text Fig. 1(e). To calibrate the vacuum optomechanical coupling rate $g_{0}$ one has to integrate both side of Eq. 1 around a mechanical sideband. Integration of $S_{\xi\xi}$ around a mechanical mode with frequency $\omega_{m}$ and OM coupling $g_{0}$ obtains $8g_{0}^{2}n_{\mathrm{th}}/k^{2}$ where $n_{\mathrm{th}}=k_{B}T/\hbar\omega$ is the thermal occupation of the mode. Using this, $g_{0}$ is given by
@@ Appendix D Frequency noise characterization
When measuring mechanical frequency fluctuations, it is known that the measurement is contaminated by three main sources of noise: thermomechanical, detection, and intrinsic frequency noise,
$\delta\omega(t)=\delta\omega_{\mathrm{th}}(t)+\delta\omega_{\mathrm{det}}(t)+\delta\omega_{\mathrm{int}}(t).$ (3)
These sources are statistically independent from each other, such that the power spectral density (PSD) of the total frequency noise is given by:
$S^{\text{tot}}_{\omega}(\Omega)=S_{\omega}^{\mathrm{th}}(\Omega)+S_{\omega}^{\mathrm{det}}(\Omega)+S_{\omega}^{\mathrm{int}}(\Omega),$ (4)
where the PSD is defined as the Fourier transform of the correlation function
$S_{\omega}^{a}(\Omega)=\int_{\mathrm{\infty}}^{\infty}\langle\delta\omega_{a}(0)\delta\omega_{a}(\tau)\rangle e^{i\omega\tau}d\tau,$ (5)
for $a=\mathrm{th},\mathrm{det},\mathrm{int}$. The detection noise originates from the noise added to the signal at the photodetection stage whereas the thermomechanical and intrinsic noise are properties of the mechanical resonator.
For the difference frequency, $\omega_{\text{d}}=\omega_{2}-\omega_{1}$, we can construct the noise model based on this model to give, {multline} δω_d(t) = δω_th,2(t) - δω_th,1(t) + δω_det,d(t)
δω_int,2(t) - δω_int,1(t). Note that while the thermomechanical and intrinsic components stem from the corresponding components in $\omega_{1}$ and $\omega_{2}$, the detection component, however, appears only once. This is because this difference frequency tone is generated within the optomechanical cavity, and prior to detection. Therefore, the contribution does not get summed in.
The corresponding frequency noise PSD is given by {multline} S_ω_d = S_ω^th,1(Ω) + S_ω^th,2(Ω)+ S_ω^det,d(Ω)
S_ω^int,1(Ω) + S_ω^int,2(Ω) - 2S_ω^int,12(Ω), where
$S_{\omega}^{\mathrm{int,12}}(\Omega)=\int_{\mathrm{\infty}}^{\infty}\langle\delta\omega_{\mathrm{int,1}}(0)\delta\omega_{\mathrm{int,2}}(\tau)\rangle e^{i\omega\tau}d\tau,$ (6)
corresponds to the correlation between the intrinsic noise components.
We fit these models to our experimental data to extract the frequency noise properties. It is known that the thermomechanical term has no frequency dependence and that the detection term has a quadratic frequency scaling. For the data shown in the main-text Fig. 1(f), we use the model $S_{\omega}(\Omega)=B\Omega^{2}+A/\Omega$ for fitting and extracting the $1/\Omega$ noise component. We have not included a constant term since, as mentioned in the main text, the thermomechanical component for the high-$Q$ OOP mode is too low to be resolved.
All frequency PSD calculations are done using the linearised LPSD python package https://pypi.org/project/lpsd/. In addition, the SciPy Cross-PSD package https://docs.scipy.org/doc/scipy/reference/generated/scipy.signal.csd.html was used to calculate frequency noise correlations.
In Fig. 6 we show an example of the full frequency noisePSD, the segment used for fitting, and the result of the fit for the lowest input optical power. The roll-off at frequencies above $1\text{\,}\mathrm{kHz}$ is a result of the finite bandwidth of the lock-in amplifier’s demodulator. We do not include this segment in the model. The initial points of the data are not included either due to the lack of sufficient number of points and large variations. $1/\Omega$-type frequency noise with a PSD $S_{\omega}(\Omega)=A/\Omega$ corresponds to a constant Allan variance with standard deviation $\sigma^{2}_{\omega}(\tau)=2A\ln{(2)}$. This relation is used to produce the inset of the main-text Fig. 1(f).
@@
Furthermore we observe and characterise the strongly correlated nature of $1/\Omega$ noise in the two modes $\omega_{1}$ and $\omega_{2}$. We do so by evaluating the frequency noise PSD (Eqn: 4) of both modes ($S_{\omega_{1}}$ and $S_{\omega_{2}}$ respectively) and then calculate the cross-PSD, given by
$S_{\omega_{a}\omega_{b}}(\Omega)=\int_{\mathrm{\infty}}^{\infty}\langle\delta\omega_{a}(0)\delta\omega_{b}(\tau)\rangle e^{i\omega\tau}d\tau,$ (7)
It is worth noting in Fig. 7 that $S_{\omega_{1}\omega_{2}}(\Omega)$ in yellow follows the same trend as that of $S_{\omega_{1}}(\Omega)$ in red. This signifies that this $1/\Omega$ contribution from both modes $\omega_{1}$ and $\omega_{2}$ is maximally correlated. As an alternate visualisation, we also plot $\bar{S}_{\omega_{1}\omega_{2}}(\Omega)=S_{\omega_{1}\omega_{2}}(\Omega)/\sqrt{S_{\omega_{1}}(\Omega)S_{\omega_{2}}(\Omega)}$, which is normalised to the point-by-point individual PSDs. It shows that the signals are correlated to nearly the same extent throughout the entire bandwidth, $\Omega$ of up to $100\text{\,}\mathrm{Hz}$.
Figure 7: frequency noise PSD of the mechanical modes $\omega_{1}$ (red) and $\omega_{2}$ (blue) as shown in main-text Fig: 2(h). Cross-PSD, $S_{\omega_{1}\omega_{2}}$ in yellow and the Correlation f-PSD normalised to the geometric mean ($\bar{S}_{\omega_{1}\omega_{2}}=S_{\omega_{1}\omega_{2}}/\sqrt{S_{\omega_{1}}S_{\omega_{2}}}$) of the individual PSDs in grey
D.1 Impact of the measurement conditions on the $1/\Omega$ noise
In the mechanical frequency fluctuations measurements, the two components that mainly affect the noise floor are measurement optical power and mechanical driving amplitude. It is expected that increasing the driving amplitude, reduces the thermomechanical and the detector contributions and increasing the optical power reduces the detector contribution as it increases the measurement SNR [42]. We verify these properties experimentally by the measuring the frequency noise PSDs for both modes for different optical powers and driving amplitudes. These measurements are shown in Fig. 8. As expected, we observe that the mechanical driving power affects both thermomechanical and detector contributions (note that for the high-Q mode the thermomechanical noise is not resolved as it is much weaker than the other contributions). We also observe that increasing the optical power consistently improves the detection noise. The other important observation is that neither optical power nor the driving power affects the $1/\Omega$ noise contribution for either of the modes. The independence of the $1/\Omega$ noise on the measurement conditions, particularly the optical power, is the first evidence for the $1/\Omega$ noise to be intrinsic to the resonator.
@@ D.1 Impact of the measurement conditions on the $1/\Omega$ noise
Appendix E Distinguishing the intrinsic $1/\Omega$ noise from environmental effects
E.1 Thermal drifts
Figure 9: Frequency noise PSD of the mechanical mode $\omega_{1}$ in comparison with the Temperature PSD, both normalised to arbitrary units. We see that the temperature noise branch has a much faster decay than the $1/\Omega$ branch at shorter time scales, implying that the processes are only correlated for timescales slower than $1\text{\times}{10}^{-2}\text{\,}\mathrm{Hz}$.
The observed frequency fluctuations at low sideband frequencies $\Omega$ have two possible causes: temperature fluctuations of the device, either due to environmental temperature drift or laser intensity fluctuations, and the intrinsic frequency flicker noise. In the following, we show that the observed frequency fluctuations can be ascribed to flicker noise alone.
Temperature variations in the environment are transduced to the polygon resonator through multiple thermal pathways, each characterized by a thermal frequency response of the generic form
$R_{\mathrm{th}}(\Omega)=(1+\Omega^{2}\tau_{\mathrm{th}}^{2})^{-1}$ (8)
where $\tau_{\mathrm{th}}$ is the characteristic thermal response time of a given body. For the chip and the assembly on which it is mounted, $\tau_{\mathrm{th}}$ is on the order of seconds [44], while for the polygon resonator, the thermal time constant is about $13.5\text{\,}\mathrm{ms}$, assuming thermal conduction occurs primarily through the tethers. Consequently, temperature fluctuations originating outside the chip are low-pass filtered by the thermal response of the chip, suppressing fluctuations faster than $\sim$1 Hz. However, in Fig. 8 we clearly observe a $1/\Omega$-power law well above 1 Hz. These frequency fluctuations cannot stem from the environment.
To visualize the difference between regimes of thermal fluctuations and flicker noise, we performed simultaneous measurements of the mechanical resonance frequency and the vacuum chamber temperature over a 7-hour period. The normalized PSDs of both quantities are shown in Fig. 9. For frequencies below $10^{-2}$ Hz, both PSDs follow a $1/\Omega$ scaling and are strongly correlated.
In order to explain the $1/\Omega$-type frequency fluctuations above 1 Hz, we consider heating sources that act locally on the device. The most likely source is intensity fluctuations of the laser light arriving on the chip. To test whether this source is responsible for the observed fluctuations, we performed frequency fluctuation measurements with different laser powers. As shown in Fig. 8, the different laser powers strongly affect the readout noise beyond 100 Hz, but leave the $1/\Omega$-type noise unchanged. For this reason, we rule out laser intensity fluctuations as the source of our frequency fluctuations between 1 and 100 Hz.
As both environmental and local temperature fluctuations are excluded as possible sources for the observed frequency fluctuations between 1 and 100 Hz, we consider intrinsic frequency flicker noise as the dominant source. Flicker noise is believed to stem from coupling of the resonator to two-level systems in the material and is expected to contribute to frequency fluctuations [15]. Clearly, we have evidence that temperature fluctuations are dominant in this regime. By contrast, above $10^{-2}$ Hz the PSD power laws decouple, with the frequency fluctuation PSD following a $1/\Omega$ law. This provides clear evidence that the mechanical frequency fluctuations at these higher frequencies arise from a non-thermal mechanism.
E.2 Gas pressure
Adsorption of gas particles on the surface of the resonator could also lead to mechanical frequency . [41]. To study the impact of this phenomenon we measure the frequency noise of the high-Q mode as a function of pressure. As shown in Fig. 10, no significant change is observed for pressures below $1\text{\times}{10}^{-4}\text{\,}\mathrm{mbar}$, implying it does not influence the intrinsic $1/\Omega$ noise. The change in the high frequency, detector-noise-limited regime is due to the reduction of the mechanical $Q$ of this mode due to gas damping and the resulting reduced oscillation amplitude.
Figure 10: Measured mechanical frequency noise PSD for $\omega_{1}$ for varying vacuum chamber pressures in the range $1\text{\times}{10}^{-6}\text{\,}$ to $1\text{\times}{10}^{-4}\text{\,}\mathrm{mbar}$
Appendix F Optical Non-linear Transduction for difference Frequency generation